# A Survey on Automatic Credibility Assessment of Textual Credibility Signals in the Era of Large Language Models

### v1

@@ 2.2. Dimensions of credibility assessment
Finally, the credibility signals differ in their polarity. In the existing works, some signals are formulated as a predictor of credible content (e.g., high factuality), while other signals are formulated as a predictor of non-credible content (e.g., presence of logical flaws). When interpreting or aggregating individual credibility signals (either by an end user or a credibility assessment algorithm), such polarity must be explicitly taken into consideration (e.g., in (Podgurski et al., 2022; Leite et al., 2024)).
@@
Importance and effects of credibility signals on end users. Due to a broad definition of credibility, as well as rich taxonomies of credibility signals/indicators, a number of human-based studies have analysed the importance of individual credibility signals and their effects on the perceived credibility, in general as well as for specific groups of end users (according to their expertise, age, education etc.).
@@
First, these studies address various content types and platforms, such as social media posts (Shusas and Forte, 2024; Al-Eidan et al., 2009), news articles (Lu et al., 2022; Huang et al., 2019; Heuer and Glassman, 2024), videos and video sharing platforms (Hughes et al., 2024; Feng et al., 2023), images (Feng et al., 2023), search engines (Wang et al., 2023; Juneja et al., 2024; Lurie and Mustafaraj, 2018), or online encyclopedias (Elmimouni et al., 2022).
@@
Regarding the scope of credibility signals, some works focus on signals provided directly by (social media) platforms (Wang et al., 2023; Lurie and Mustafaraj, 2018), some investigate a single credibility signal (e.g., hyper-partisanship in (Jia et al., 2022)) or a whole spectrum of signals (e.g., 28 signals analysed in (Piccolo et al., 2021)). Jia et al. (2022) compared credibility signals produced by three various sources – an algorithm, a community, and a third-party fact-checker. Besides content- and context-based signals, Chang et al. (2021) considered also an additional group of design signals (consisting of four signals: interaction design, interface design, navigation design, and security settings).
The number of studies comparing the importance of credibility signals on non-English data is rather limited. The work by Gao et al. (2015) performs the human study on the importance of credibility signals for the Weibo Chinese misinformation dataset regarding health and safety information.
@@ 2.3. Existing surveys
The majority of the surveys on information credibility focus on specialised domains. Alrubaian et al. (2019) conducted a literature review on the credibility of social media information while considering 4 levels of credibility assessment (i.e., post, member, topic/event, hybrid level). The analysed credibility analysis methods were categorised into automated-, human-, and hybrid-based approaches. Automated-based approaches are further subcategorized into machine learning approaches, graph-based semisupervised approaches, and weighted and IR algorithms. Their analysis spans until 2018, and therefore, the automated methods are limited to traditional machine learning (e.g., classifications algorithms like decision trees, logistic regression) and do not include more recent transformer or LLM-based approaches.
Similarly, Qureshi et al. (2021) perform an analysis of the credibility detection methods in the social media and microblog domains. The authors adopt the level-based classification of the task proposed by Alrubaian et al. (2019). In addition, they propose a fine-grained framework based on the existing approaches that targets various aspects behind credibility assessment. Their theoretical framework consists of content-, media-, website- and interaction-based assessment of social media information. Within this framework, they identify two main directions of research, namely, post-level and user-level constructs. The post-level constructs are further divided into the detection of: (i) deceptive and trustful information (rumor and fake news, hoax, spam and phishing and meme detection tasks), (ii) bias and objectivity detection (hyperpartisan, bias and polarisation detection tasks), and (iii) hate speech and offensive language detection tasks. The user-based constructs, on the other hand, are based on: (i) user deception level (bot and suspicious behaviour detection tasks), (ii) user competence (topic-specific expert detection; and opinion credibility detection tasks), and (iii) user authority (post ranking, personality detection, influence and trust propagation).
Another group of the surveys focuses on particular categories of credibility signals, such as, a survey by Nakov et al. (2021b) that surveys the literature on factuality and biases in social media, or a survey by Panchendrarajan and Zubiaga (Panchendrarajan and Zubiaga, 2024) on claim detection for automated fact checking.
@@ 2.4. Scope of the survey
Following the existing surveys and recent development in the relevant research field, we define the scope of this survey as follows:
Out of approaches to study credibility, this survey specifically tackles automation of credibility assessment.
Regarding the levels of credibility assessment, the research works falling into the scope of this survey typically address the credibility assessment at the post level, but indirectly also at user/source, topic/event, media and hybrid levels.
Out of various taxonomies of credibility signals, the survey covers content- as well as context-based signals that relate to text (while relation to other modalities is briefly documented in Section 8.7). We specifically focus on a selection of such credibility signals that: (i) have attracted a considerable amount of interest in the NLP community, and (ii) have been recognized by the human-based studies as important and impactful on social media users and human experts (e.g., media professionals). Firstly, the survey covers papers that address automatic credibility assessment with multiple categories of credibility signals at once. Subsequently, it provides an in-depth analysis for three individual categories of credibility signals:
@@
Claims and veracity.
For survey completeness, we provide a brief analysis for six additional categories of credibility signals, which are either covered by recent surveys (e.g., machine-generate text, offensive language) or have been addressed by a limited number of works yet:
Machine-generated text;
Text quality;
References and citations;
Clickbaits and title representativeness;
Originality and content reuse;
Offensive language.
@@ 3.1. Research paper collection process
Credibility signal category | Description | Search query | Search scope | ||||||||||||||||
@@
Title, abstract, keywords | ||||||||||||||||
Title, abstract, keywords | ||||||||||||||||
Claims and veracity |
Title, abstract, keywords | ||||||||||||||||
@@
Identification – Search queries. In order to conduct a systematic literature search that covers the papers that have a high potential of being relevant, we generated custom search queries for papers tackling with: (i) credibility assessment utilizing multiple categories of credibility signals, as well as (ii) automatic detection of individual categories of credibility signals. Each query combined keywords and logical search operations (AND, OR) to achieve the highest possible coverage and reduce false positives at the same time. The keywords have been carefully selected to capture inconsistent and ambiguous terminology used in the existing works. The first version of queries was proposed following our past experience and research in the respective area, and iteratively refined to cover additional alternative phrases with the same semantics identified in the research papers. Table 1 represents the search queries employed to acquire the desired list of research publications.
Identification – Databases and registers. The proposed queries were submitted to four large repositories of peer-reviewed scientific publications, namely:
ACL Anthology444https://aclanthology.org/
ScienceDirect555https://www.sciencedirect.com/
ACM DL666https://dl.acm.org/
IEEE Xplore777https://ieeexplore.ieee.org/
Identification – Search scope. Papers addressing credibility assessment with multiple categories of credibility signals are quite scarce, therefore, for this type of papers we performed the search for the query terms in the whole body of the document. For individual categories of credibility signals, on the other hand, we searched for the given query term in the title, abstract and user-specified keywords888The search in author-specified keywords is available only at the ScienceDirect digital library. only.
@@
Screening. Upon identification of the papers that satisfy the search queries (n = 13,243), title and abstract of the papers have been screened directly in the user interface of corresponding digital library. Following this screening, the non-relevant papers (n = 13,010) have been excluded. As excluded papers in this step, we considered those that: (i) used the search keywords in a different context that is out of scope this survey as defined in Section 2.4 (i.e., not related to online content credibility and NLP), or (ii) are not full/short/journal/workshop papers (e.g., abstracts of invited talks or workshops).
Out of the remaining relevant papers (n = 233), the full body (PDF file) cannot be retrieved by the authors of this survey only in a negligible number of cases (n = 9).
For all papers with a full body (n = 224), a final eligibility check was done (to further exclude such papers that have not been already excluded in the first screening step). By in-depth analysis of the full version of the obtained papers, we excluded papers due to two specific reasons:
Non-NLP (n = 38 for direct search, n = 6 for extended search) – the paper is exploring the general or specific notion of credibility and credibility signals, but from either theoretical, behavioural, human annotation, or other area of computer science (computer vision, social network analysis, etc.).
Non-relevant (n = 53 for direct search, n = 5 for extended search) – the search terms (co)-occurr in the paper in the context different from the notion of credibility signals. For example, the bi-gram “confidence indicator” can often occur in research on construction risk assessment. Such papers were removed from the analysis.
Included. Finally, we obtained relevant NLP papers (n = 175) that are exploring a specific category or multiple categories of credibility signals while the primary contribution is related to NLP domain (i.e., a paper performs an automatic detection and/or aggregation of credibility signals, introduces a novel dataset or a system for automatic credibility assessment).
Out of included papers, the smallest number of papers tackles with credibility assessment with multiple categories of credibility signals, while the most populated is claims and veracity (see Figure 3). Due to NLP focus, the ACL Anthology is clearly the most common digital library for all three specific categories of credibility signals, while number of papers from individual digital libraries is rather evenly distributed for multi-category credibility assessment papers (see Figure 3). From the time evolution, we can see an initial increase of papers in 2016, corresponding to the overall research focus on false information after the US presidential elections (see Figure 4). A significant increase in recent years, partially slow downed during the COVID-19 pandemic, demonstrates a high research interest in the area addressed in this survey.
@@ 3.2. Paper categorization and in-depth analysis
human agreement (if paper introduces an annotated dataset, what kind of human agreement was achieved).
The full list of papers included in this survey, including all annotations is available as a digital appendix of this survey999Available at: https://docs.google.com/spreadsheets/d/1-6ueO3a_VxgBuiv0QluUY3Un9YEps4YPhtbtESdTaTM/edit?usp=sharing.
Following this analysis, we describe multi-category credibility assessment papers as well as each individual category of credibility signals from three perspectives (that are also utilized to structure the consequent categories): (i) datasets, (ii) methods and models, and (iii) tools and services. For each category we summarized the current state of the art and identified problems and challenges that are specifically present in such credibility signal category (such analysis is further elaborated in Section 9, with an orthogonal discussion across all included papers).
4. Credibility assessment with multiple categories of credibility signals
This section provides an overview of research works that tackle with credibility assessment by considering multiple categories of credibility signals at once. This group of approaches typically explicitly uses the terminology relevant to credibility (such as credibility assessment, credibility signal, credibility score), in contrast to approaches described in the following sections that tackle with detection of a particular category of credibility signals, but typically do not denote them as such.

### v2

@@ 2.2. Dimensions of credibility assessment
Second, the credibility signals differ in their polarity. In the existing works, some signals are formulated as predictors of high credibility (e.g., high objectivity), while other signals are formulated as predictors of low credibility (e.g., presence of logical flaws). When interpreting or aggregating individual credibility signals (either by an end user or a credibility assessment algorithm), such polarity must be explicitly taken into consideration (e.g., in (Podgurski et al., 2022; Leite et al., 2024)).
Third, Credible Web Community Group (Multiple authors and Hawke, 2019) structured credibility signals by a subject type (a part of target entity they are derived from): claim, text, image, audio, video, article, title, web page, website, aggregation (e.g., Really Simple Syndication - RSS), venue, provider, creator, person, or organization.
Despite many commendable efforts, there is no standardized unified list and categorization of credibility signals. As a result, research works commonly utilize various and very diverse lists of credibility signals. For example, Molina et al. (2021) proposed separate lists of credibility signals (or features as they have been denoted by the authors) to distinguish between eight types of online content, such as real news, fabricated news, or satire. Each list grouped signals into 4 main categories roughly corresponding to levels of credibility assessment: (i) message and linguistic, (ii) sources and intentions, (ii) structural, and (iv) network.
An alternative approach was used by the Credibility Coalition222https://credibilitycoalition.org/, which organized weekly remote sessions during which participants drafted about 100 credibility signals specifically aimed at the credibility of web pages (Zhang et al., 2018). They have been later coalesced into 12 major categories, including reader behaviour, revenue models, publication metadata, and inbound and outbound references. In 2017, the Credibility Coalition formed the World Wide Web Consortium (W3C) Credible Web Community Group, which continued in this initiative. It crowdsourced from human experts (researchers as well practitioners) an extensive list containing more than 200 credibility signals, commonly referred to as W3C signals (Multiple authors and Hawke, 2019). This list introduces a number of signals categorized by subject types. While such a list still remains an informal and incomplete draft and has not been standardized yet, it has been adopted and served as a foundation to several research works (e.g., (Piccolo et al., 2021; Podgurski et al., 2022)).
We can conclude that no standard or widely used and recognized list or categorization of credibility signals currently exists that could serve as a solid basis for our literature survey. Categorization by means of subject types as proposed by W3C signals (Multiple authors and Hawke, 2019) is probably the most comprehensive and elaborated one, nevertheless, it mixes at the top level different modalities (text, video, audio, etc.) and content types (article, title, claim etc.). More specifically, the signals relevant to our survey are not optimally classified under the text modality or across various individual content types. Therefore, we thoroughly analysed the relevant existing works (Molina et al., 2021; Zhang et al., 2018; Multiple authors and Hawke, 2019; Podgurski et al., 2022; Nakov et al., 2024) and proposed a novel unified categorization of textual credibility signals (see Figure 2). We adopted a bottom-up approach as follows:
First, out of mentioned credibility signals, we identified specifically textual credibility signals (i.e., signals that can be derived from textual content by means of diverse NLP methods).
We grouped the resulting credibility signals according to their semantic similarity and shared datasets and methods applied for their detection. To this end, we employed three grouping strategies. At first, we considered the existing categorizations that have been proposed by the relevant works, especially individual (sub)categories from the above-mentioned W3C signals (Multiple authors and Hawke, 2019). In this way, we grouped all signals related to a title into a single category named Clickbaits and title representativeness, all signals related to claims into a category of Check-worthy and fact-checked claims as well as reused a subcategory of References and citations. Second, we grouped together signals which refer to the same or highly-similar concepts and just use a different terminology, e.g., grammar and spelling errors, grammar, spelling or punctuation mistakes, punctuation naturally fall under the category of Text quality; or profanity, incivility, impoliteness and hate speech under the category of Offensive language. In some cases, the signals refer to the same concept, but in an opposite meaning, like originality and attribution of non-original content belong to the same category of Originality and content reuse. In some cases, the signals refer to the same concepts but are defined on different granularity levels, e.g., logical fallacies, logical flaws, ad-hominem attacks all together belong under Persuasion techniques and logical fallacies. Third, we grouped together signals where the borders between them are not clearly defined and in practice they correlate with each other, e.g., subjectivity, political bias, confirmation bias, emotional tone were grouped under Factuality, subjectivity and bias. We also jointly combined these strategies, especially the category of Factuality, subjectivity and bias groups signals that are closely related, commonly refer to the same concepts but from opposite angles, or are at different levels of abstractions.
In this way, we finally proposed 8 categories of credibility signals (see also Figure 2; more detailed description of each category is provided in the later sections): (i) Factuality, subjectivity and bias; (ii) Persuasion techniques and logical fallacies; (iii) Check-worthy and fact-checked claims; (iv) Text quality; (v) References and citations; (vi) Clickbaits and title representativeness; (vii) Originality and content reuse; and (viii) Offensive language.
In addition, we recognized one more category of signals – (ix) Machine-generated text – that has become very relevant only recently with the emergence of LLMs and, therefore, it did not appear in the existing lists/taxonomies. At the same time, it is a strong predictor of credibility as fully generated text can contain many factual errors either as a result of unintended hallucinations of LLMs or due to their intended misuse to generate false content (the previous works demonstrated significant vulnerabilities of LLMs to generate disinformation (Vykopal et al., 2024), even personalized one (Zugecova et al., 2025)).
@@
Importance and effects of credibility signals on end users. Due to a broad definition of credibility, as well as rich taxonomies of credibility signals, a number of human-based studies have analysed the importance of individual credibility signals and their effects on the perceived credibility, in general as well as for specific groups of end users (according to their expertise, age, education etc.).
@@
First, these studies address various content types and platforms, such as social media posts (Shusas and Forte, 2024; Al-Eidan et al., 2009), news articles (Lu et al., 2022; Huang et al., 2019; Heuer and Glassman, 2024), videos and video sharing platforms (Hughes et al., 2024; Feng et al., 2023), images (Feng et al., 2023), search engines (Wang et al., 2023a; Juneja et al., 2024; Lurie and Mustafaraj, 2018), or online encyclopedias (Elmimouni et al., 2022).
@@
Regarding the scope of credibility signals, some works focus on signals provided directly by (social media) platforms (Wang et al., 2023a; Lurie and Mustafaraj, 2018), some investigate a single credibility signal (e.g., hyper-partisanship in (Jia et al., 2022)) or a whole spectrum of signals (e.g., 28 signals analysed in (Piccolo et al., 2021)). Jia et al. (2022) compared credibility signals produced by three various sources – an algorithm, a community, and a third-party fact-checker. Besides content- and context-based signals, Chang et al. (2021) considered also an additional group of design signals (consisting of four signals: interaction design, interface design, navigation design, and security settings). While the most of these studies focused on English content, the number of studies comparing the importance of credibility signals on non-English data is rather limited. The work by Gao et al. (2015) performs the human study on the importance of credibility signals for the Weibo Chinese misinformation dataset regarding health and safety information.
@@ 2.3. Existing surveys
The existing surveys on automatic credibility assessment focus on specialised domains. Alrubaian et al. (2019) conducted a literature review on the credibility of social media information while considering 4 levels of credibility assessment (i.e., post, member, topic/event, hybrid level). The surveyed methods were categorised into automated-, human-, and hybrid-based approaches. Automated-based approaches are further subcategorised into machine learning approaches, graph-based semisupervised approaches, and weighted and IR algorithms. Their analysis spans until 2018, and therefore, the automated methods are limited to traditional machine learning (e.g., classifications algorithms like decision trees, logistic regression) and do not include more recent transformer or LLM-based approaches. Similarly, Qureshi et al. (2021) perform an analysis of the credibility detection methods in the social media and microblog domains. The authors adopt the level-based classification of the task proposed by Alrubaian et al. (2019). In addition, they propose a fine-grained framework based on the existing approaches that targets various aspects behind credibility assessment. Their theoretical framework consists of content-, media-, website- and interaction-based assessment of social media information. Within this framework, they identify two main directions of research, namely, post-level and user-level constructs. The post-level constructs are further divided into the detection of: (i) deceptive and trustful information, (ii) bias and objectivity detection, and (iii) hate speech and offensive language detection tasks. The user-based constructs, on the other hand, are based on: (i) user deception level, (ii) user competence, and (iii) user authority.
Another group of the surveys focuses on particular categories of credibility signals. For example, Nakov et al. (2024) surveyed the literature on factual accuracy and biases in social media, Panchendrarajan and Zubiaga (2024) analysed claim detection for automated fact checking, or Wu et al. (2025) addressed machine-generated text detection.
@@ 2.4. Scope of the survey
At first, this survey aims to systematically cover research that explicitly addresses the entire process of automatic credibility assessment (i.e., both steps of credibility signals detection as well as their aggregation into a single credibility label/score). However, restricting our review solely to such works would overlook a substantial body of research focused on the detection of individual categories of credibility signals. Although these studies may not explicitly frame their contributions within the context of credibility assessment, i.e., they often do not adopt the relevant credibility-related terminology or perform aggregation of multiple signals, they provide valuable insights and methods that are highly relevant to the broader credibility assessment framework.
More specifically, we define the scope of this survey as follows:
Out of methods to credibility assessment, this survey specifically tackles automation-based approaches.
Regarding the levels of credibility assessment, the research works falling into the scope of this survey typically address the credibility assessment at the content level, but indirectly also at user/source, topic/event, media and hybrid levels.
Out of various taxonomies of credibility signals, the survey covers both content- as well as context-based signals, predicting high as well as low credibility that are text-based (while relation to other modalities is briefly documented in Section 8.7).
Regarding the categorization of credibility signals, the survey systematically covers all 9 categories of textual credibility signals from the unified categorization of textual credibility signals (as shown in Figure 2).
Out of 9 categories of textual credibility signals, we put a specific focus and provide an in-depth analysis for three of them, namely:
@@
Check-worthy and fact-checked claims.
This selection is motivated by several factors. (i) The selected categories have attracted a considerable amount of interest in the NLP community, resulting in a substantial body of work that is suitable for systematic review. Moreover, the authors of this survey have direct practical experience in training and deploying detection models for these specific signals, which enables a more informed and application-oriented analysis and discussion. (ii) Although some of these categories have been partially addressed in the previous surveys, they have not been examined through the lens of credibility and credibility assessment, nor in terms of their integration with other signal detection approaches. Importantly, these categories offer strong potential for effective methodological integration. For example, subjectivity, bias, and persuasion techniques share underlying linguistic features, which can be exploited through transfer learning or multitask learning applied to shared datasets. (iii) The selected categories have been also identified in human-centered studies as particularly influential in shaping perceptions among social media users and domain experts (e.g., journalists and media professionals) (see, e.g., Piccolo et al. (2021) and Chang et al. (2021), and Section 2.2 for further discussion).
@@ 3.1. Research paper collection process
| Survey section | Description | Search query | Search scope | ||||||||||||||||||
@@
Title, abstract, keywords | ||||||||||||||||||
Title, abstract, keywords | ||||||||||||||||||
Title, abstract, keywords | ||||||||||||||||||
@@
Identification – Search queries. To conduct a systematic literature search aimed at identifying papers with high potential relevance, we formulated custom search queries (see Table 1) targeting two main parts of this survey: (i) studies that explicitly address the credibility assessment process, including both the detection of credibility signals and their aggregation, and (ii) studies focused on the automatic detection of the three selected individual categories of credibility signals. Each query combined keywords and logical search operations (AND, OR) to achieve the highest possible coverage and reduce false positives at the same time. The keywords have been carefully selected to capture inconsistent and ambiguous terminology used in the existing works. The first version of queries was proposed following our past experience and research in the respective area, and iteratively refined to cover additional alternative phrases with the same semantics identified in the research papers. If a search query returned a paper that was more appropriately aligned with a different category of credibility signals, the paper was reassigned accordingly. Thanks to the use of mutually distinctive categories of credibility signals, no duplicates were present across the identified sets of papers. Additional categories of credibility signals were covered through ad hoc searches, conducted without predefined search queries.
Identification – Databases and registers. The proposed queries were submitted to four large repositories of peer-reviewed scientific publications, namely: (i) ACL Anthology333https://aclanthology.org/; (ii) ScienceDirect444https://www.sciencedirect.com/; (iii) ACM DL555https://dl.acm.org/; and (iv) IEEE Xplore666https://ieeexplore.ieee.org/.
Identification – Search scope. Papers explicitly addressing credibility assessment and credibility signals detection are quite scarce, therefore, for this type of papers we performed the search for the query terms in the whole body of the document. For individual categories of credibility signals, on the other hand, we searched for the given query term in the title, abstract and user-specified keywords777The search in author-specified keywords is available only at the ScienceDirect digital library. only.
@@
Screening. After identifying all papers matching the search queries (n = 13,243), we performed an initial screening of titles and abstracts directly within the user interfaces of the respective digital libraries. Based on this step, we excluded non-relevant papers (n = 13,010). Papers were excluded at this stage if they: (i) used the search keywords in a context that falls outside the scope of this survey as defined in Section 2.4 (e.g., not related to online content credibility or NLP), or (ii) were not full, short, journal, or workshop papers (e.g., abstracts of invited talks or workshops).
From the remaining pool of potentially relevant papers (n = 233), the full-text PDF could not be retrieved for only a negligible number (n = 9).
For all papers with an accessible full text (n = 224), a final eligibility assessment was performed to exclude any remaining irrelevant papers not identified during the initial screening. Based on an in-depth analysis of the full papers, exclusions were made for the following two reasons:
Non-NLP (n = 38 for direct search, n = 6 for extended search): These papers explored credibility or credibility signals from theoretical, behavioral, human annotation, or other non-NLP perspectives (e.g., computer vision, social network analysis).
Non-relevant (n = 52 for direct search, n = 5 for extended search): These papers mentioned the search terms but in contexts unrelated to credibility signals (e.g., “confidence indicator” in construction risk assessment). Such papers were excluded from further analysis.
Included. Ultimately, we identified a set of relevant NLP papers (n = 175) that addressed either credibility assessment or one of selected credibility signal categories, with a primary focus on the NLP domain (i.e., papers that proposed automatic detection or aggregation methods for credibility signals, introduced relevant datasets, or developed systems for automatic credibility assessment).
Among the included works, the smallest group consists of papers that explicitly address credibility assessment, whereas the most populated category is check-worthy and fact-checked claims (see Figure 5). Due to the survey’s focus on NLP, the ACL Anthology emerges as the most prominent source for all three individual categories of credibility signals. In contrast, papers addressing credibility assessment are more evenly distributed across digital libraries (see Figure 5).
In terms of publication trends, we observe an increase beginning in 2016, likely corresponding to heightened research attention following the U.S. presidential elections (see Figure 6). The continued growth in subsequent years – despite a temporary slowdown during the COVID-19 pandemic – demonstrates sustained and growing interest in the topic addressed by this survey.
@@ 3.2. Paper categorization and in-depth analysis
human agreement (if paper introduces an annotated dataset, what kind of annotator mutual agreement was achieved, measured by standard metrics, e.g., by Cohen’s Kappa).
The full list of papers included in this survey, including all annotations is available as a supplementary material to this article in the ACM Digital Library888A full list of papers is available as supplementary material to this article in the ACM Digital Library and at https://kinit.sk/public/acm-tist-credibility-assessment-survey.html.
Following this analysis, we describe credibility assessment papers as well as each individual category of credibility signals from three perspectives (that are also utilized to structure the consequent categories): (i) datasets, (ii) methods and models, and (iii) tools and services. For each category we summarised the current state of the art and identified problems and challenges that are specifically present in such credibility signal category (such analysis is further elaborated in Section 9, with an orthogonal discussion across all included papers).
4. Automatic credibility assessment
This section provides an overview of research efforts that address automatic credibility assessment, encompassing both steps of the process: the detection and subsequent aggregation of credibility signals. In some cases, however, the features used as input to the credibility assessment algorithms are relatively shallow, relying primarily on surface-level linguistic characteristics (e.g., the number of unique words), which only approximate underlying credibility signals rather than explicitly capturing them. This group of approaches explicitly uses the terminology relevant to credibility (such as credibility assessment, credibility signal, credibility score), in contrast to approaches described in the following sections that address detection of a particular category of credibility signals, but typically do not denote them as such.